Q No.-1> atoms may be regarded as comprising of p's,n's and e's.if the mass attributed by electrons was doubled and that by neutrons was halved d at mass of C-12 'll b?


a)approx. the same


b)doubled


c)reduced aprrox. 25%


d)approx. halved


Q No 2> the wave no. fo 1st line of Balmer series of H z 15200cm^-1.the wave no of the first balmer line of Li2+ ion z?


a)15200


b)60800


c)76000


d)136800

Grade:12

2 Answers

Ramesh V
14 years ago

(1)

C-reduced approx. by 25 %

(2)

For atoms ions ,

                           R*(1/n2)*Z2

For Lithium , its    = 15000*32

                         = 136800

--

regards

Ramesh

deeksha sharma
13 years ago

1) c reduced by 25%.

reason --- electrons do not constitute the mass. so their doubling would bring negligible effect. halving of neutrons. will let them 3

thus the mass is p+n   6+3=9 i.e. reducedby 25%

2) FOR H like atoms, it is n2*wave no for H

=15200*3=136800